OpenStudy (anonymous):

Please help! Step by step?? 1. 5^x = 4^x+3 and 2. .16^4+3x = .3^8-x how do i solve these? step by step please? the answers are supposed to be: 1. 3log4 / log 1.25 2. (log .3^8 / .16^4 ) / (log (.16^3 x .3) help?

OpenStudy (phi):

as a first step I would use the rule \[ \log( a^b) = b \log(a) \] take the log of both sides of the equation, and then use that rule. what do you get?

OpenStudy (anonymous):

X log 5 = X+3 log 4

OpenStudy (phi):

yes but put parens around the x+3 like this x log 5 = (x+3) log4 remember that log4 is just a number (use a calculator to find it... if you are curious what number). like any number, you can "distribute" it which means multiply it times everything inside the parens. you get x log 5 = x log 4 + 3 log 4 subtract x log 4 from both sides: x log 5 - x log 4 = x log 4 - x log 4 + 3 log 4 of course x log 4 - x log 4= 0 so this is the same as x log 5 - x log 4= 3 log 4 factor x out of each term on the left side: x (log 5 - log 4) = 3 log4 can you finish ?